OpenStudy (anonymous):

Do the sides lengths 13, 84, and 85 form a Pythagorean Triple? Why or why not?

OpenStudy (btaylor):

yes, because 13^2 + 84^2 = 85^2

jimthompson5910 (jim_thompson5910):

Use the pythagorean theorem to check a^2 + b^2 = c^2 13^2 + 84^2 = 85^2 169 + 7056 = 7225 7225 = 7225 Since that last equation is true, this means that these side lengths do form a right triangle

jimthompson5910 (jim_thompson5910):

So 13, 84, and 85 form a Pythagorean Triple.